a bag has 9 tickets marked with the number 1,2,3,4,5,6,7,8,9.two tickets are drawn at random from the bag.find the probability that both numbers drawn are
A)even
B)odd

Answer by stanbon(75887)

a bag has 9 tickets marked with the number 1,2,3,4,5,6,7,8,9.two tickets are drawn at random from the bag.find the probability that both numbers drawn are
A)even
There are 4 even digits.
P(even) = 4C2/9C2 = 6/36 = 1/6
--------
B)odd
There are 5 odd digits.
P(odd) = 5C2/9C2 = 10/36 = 5/18
=============================
Cheers,
Stan H.